OpenStudy (anonymous):

Determine the sum of the measures of the interior angles of the specified polygon. A hexagon A. 900° B. 1080° C. 720° D. 540°

OpenStudy (anonymous):

Exterior = 360/6 = 60 so Interior = 120 and 120*6 = 720